OpenStudy (anonymous):

I'll become your fan and give you a medal if you get this right! Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2.4 hours, and Car B traveled the distance in 4 hours. Car A traveled 22 miles per hour faster than Car B. How fast did Car A travel?

OpenStudy (shinalcantara):

since both cars covered the same distance then Velocity A=x/2.4 and Velocity B=x/4 Velocity A=x/4 + 22 as it travels 22mph faster than B. Now try to equate the two Velocity A having different expressions.

OpenStudy (shinalcantara):

what you've got?

OpenStudy (shinalcantara):

If you'll be able to get the value of x, then divide it with 2.4 hrs. then you'll have the velocity of car A